# Toward a unified theory for common affine roots of general sets of multivariate polynomials

### v2

@@ 2 Main results on multivariate polynomials
originally introduced for finite fields, but here applied to the case of any field. Clearly, for general $F_{1},\ldots,F_{k}$ as in Corollary 10 the smallest attainable number of non-common roots in $A_{1}\times\cdots\times A_{m}$ equals $M_{k}(C_{1},C_{2})$. In the coming sections we shall demonstrate that if $A\subseteq A_{1}\times\cdots\times A_{m}$ satisfies $A<M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ then there exist $k$ linearly independent polynomials as in Theorem 5 which all evaluates to 0 on $A$, but that for some $A\subseteq A_{1}\times\cdots\times A_{m}$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ this does not hold. Here, the dual space (or null-space) is with respect to the usual inner product. Finally, as we shall demonstrate in the next two sections of the paper the right hand side of (1) serves as an upper bound on $M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ and we already know that $M_{k}(C_{1},C_{2})$ equals the right hand side of (3). We conclude that problem (I) and problem (II) are basically identical, but applied to dual spaces. All claims of the present remark follows by inspection of the proof of Theorem 5 in Section 4.
We conclude this section by demonstrating the superiority of the footprint bound over other known bounds on roots of general single mulitvariate polynomials. Among these the generalized Alon-Füredi bound [7] is the sharpest as it has as corollary the Alon-Füredi bound [2] which in turn has as corollaries [7] Alon’s combinatorial Nullstellensatz [1] as well as the Schwartz-Zippel bound [42, 36]. For common roots of sets of more general multivariate polynomials we are not aware of other efficient bounds than the footprint bound.
By inspection the generalized Alon-Füredi bound [7, Th. 1.2] can be reformulated as follows.
@@
In [6] it is already noted that the generalized Alon-Füredi bound is a consequence of the footprint bound. We here give some more details including discussing when the two bounds may or may not coincide. Consider a graded monomial ordering, i.e. a monomial ordering $\prec$ satisfying that if $\deg N_{1}<\deg N_{2}$ then necessarily $N_{1}\prec N_{2}$. There are many such orderings, including $m!$ degree lexicographic ordering where in addition to the above property ties are broken by the rule that if $K_{1}$ and $K_{2}$ are of the same degree then $K_{1}\prec K_{2}$ if $K_{1}\prec_{lex}K_{2}$ for the given chosen lexicographic ordering. For a degree monomial ordering for sure the leading monomial ${\mbox{lm}}(F)$ belongs to the set considered in (5) and Theorem 12 follows imediately from Corollary 10.
We now comment on the possible tightness of the generalized Alon-Füredi bound. If $d_{1}+\cdots+d_{m}=d$ then there is only one monomial in the support ${\mbox{Supp}}(F)$ of highest degree, hence the two bounds coincide. Moreover, $F$ is a monomial ordering invariant polynomial [21, Def. 8], i.e. all monomials in the support of $F$ divides the unique monomial of highest degree. As shown in [21, Cor. 9] the only polynomials that attain the footprint bound in such a case are the products of linear terms as described in the proof of Corollary 10 (see also [7, Sec. 3.3]). Imagine next that $d_{1}+\cdots+d_{m}\neq d$, in which case there need to be more than one monomial in the argument of (5). If for some $i$ no monomial $M$ in the support of $F$ satisfies simultaneously $\deg_{X_{i}}M=d_{i}$ and $\deg M=d$ then for a lexicographic ordering with $X_{j}\prec_{lex}X_{i}$ for all $j\neq i$ the footprint bound will produce a strictly tighter result than Theorem 12. Example 2 below illustrates the situation. Assume on the other hand that for all $i$ such a monomial exists, then a nesecary condition for Theorem 12 to produce as sharp a result as the footprint bound is that for each of the degree lexicographic orderings the largest monomial in the argument of (5) is of $\sigma$-value equal to (5) which indeed is a very restrictive assumption on the form of $F$.
@@ 3 Results in linear algebra setting
As a preperation for providing in Section 4 a proof of Theorem 5 as well as giving the missing details of Remark 11 in the current section we treat problems similar to (I) and (II), but in a much more general context than that of multivariate polynomials. Admittedly, for that purpose it would be enough to consider in this general setting only problem (II), as we already proved Theorem 8, Corollary 9 and Corollary 10, but for completeness and to support possible future applications we shall also treat (I).
@@
In this section we start by enhancing such results to hold over any field $\mathbb{F}$}, i.e. we demonstrate that the proof of the Feng-Rao bounds hold regardless of underlying field. Our exposition closely follows that of [3, 25]. With such theory in hand and by applying Forney’s two duality lemmas [19] our results concerning multivariate polynomials follows as easy corollaries, but admittedly our exposition in the present section is more general than what is needed for that purpose as we want to support possible future work on generalizing our results in Section 2 to also hold over other algebraic structures than fields.
@@
${\mbox{Supp}}(D)=\{i\in{\mathcal{I}}\mid c_{i}\neq 0{\mbox{ for some }}(c_{1},\ldots,c_{n})\in D\}.$
We are interested in establishing information on the size of ${\mbox{Supp}}(D)$ which we call $w_{H}(D)=\#{\mbox{Supp}}(D)$. Here, the subscript $H$ refers to the Hamming weight from coding theory.
Let $B=\{{\vec{b}}_{1},\ldots,{\vec{b}}_{n}\}$ be an ordered basis for ${\mathbb{F}}^{n}$. Define $L_{0}=\{\vec{0}\}$ and for $i=1,\ldots,n$ $L_{i}:={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{i}\}$. Further define $\bar{\rho}:{\mathbb{F}}^{n}\rightarrow\{0,1,\ldots,n\}$ by $\bar{\rho}(\vec{0})=0$ and for $\vec{c}\neq\vec{0}$ $\bar{\rho}(\vec{c})=i$ if $\vec{c}\in L_{i}\backslash L_{i-1}$ and $m(\vec{c})=\min\{m\mid\vec{c}\cdot\vec{b}_{m}\neq 0\}$. Here, $\cdot$ is the usual inner product.
@@
Consider a subspace $D\subseteq{\mathbb{F}}^{n}$ of dimension $k\geq 1$. There exists a basis $\{\vec{c}_{1},\ldots,\vec{c}_{k}\}$ for $D$ with $\bar{\rho}(\vec{c}_{1}),\ldots,\bar{\rho}(\vec{c}_{k})$ being pairwise different. For such a basis it holds that $m(\vec{c}_{1}),\ldots,m(\vec{c}_{k})$ are also pairwise different. Both the $\bar{\rho}$-values and the $m$-values are invariants of $D$. If $\vec{c}\in D\backslash\{\vec{0}\}$ then $\bar{\rho}(\vec{c})$ and $m(\vec{c})$, respectively, belongs to the above mentioned sets, respectively.
Hence, the following definition is clear
$\bar{\rho}(D)=\{e_{1},\ldots,e_{k}\mid 1\leq e_{1}<\cdots<e_{k}\leq n,\forall e_{i}\exists\vec{c}\in D\backslash\{\vec{0}\}{\mbox{ with }}\vec{c}\in L_{e_{i}}\backslash L_{e_{i}-1}\}$
$m(D)=\{m_{1},\ldots,m_{k}\mid 1\leq m_{1}<\cdots<m_{k}\leq n,\forall m_{i}\exists\vec{c}\in D\backslash\{\vec{0}\}{\mbox{ with }}\vec{c}\in L^{\perp}_{m_{i}-1}\backslash L^{\perp}_{m_{i}}\}$
We next introduce the concept of one-way well-behaving pairs which is more general than the more well-known concepts of well-behaving pairs and weakly well-behaving pairs [32] allowing for a more powerful description [25]. Consider an additional ordered basis $B^{\prime}=\{\vec{b}^{\prime}_{1},\ldots\vec{b}^{\prime}_{n}\}$ (which may or may not be equal to $B$). Let ${\mathcal{I}}:=\{1,2,\ldots,n\}$. An ordered pair $(i,j)\in{\mathcal{I}}^{2}$ is said to be one-way well-behaving (OWB) if $\bar{\rho}(\vec{b}_{u}\ast\vec{b}^{\prime}_{j})<\bar{\rho}(\vec{b}_{i}\ast\vec{b}^{\prime}_{j})$ for $u<i$. Given bases $B,B^{\prime}$ as above consider for $l=1,2,\ldots,n$ and $i=1,2,\ldots,n$ the following sets
@@
To prove (10) we therefore repeat the proof of [3][Th. 10] which with our notation goes as follows. Write $e_{u}:=\bar{\rho}(\vec{c}_{u})$ for $u=1,\ldots k$ and observe that if $(e_{u},j)$ is OWB for some $j\in\{1,\ldots,n\}$ and $\bar{\rho}(\vec{b}_{e_{u}}\ast\vec{b}^{\prime}_{j})=\ell$ then by the very definition of OWB also $\bar{\rho}(\vec{c}_{u}\ast\vec{b}_{j}^{\prime})=\ell$. Hence,
@@
To prove (11) we repeat the proof of the last part of [25][Th. 1]. Write $\gamma:=\bar{\mu}(m_{1}:=m(\vec{c}_{1}),\ldots,m_{k}:=m(\vec{c}_{t}))$ and
@@
For the purpose of the present paper the reader may think of $\mathcal{B}$} and $\mathcal{B}$$^{\prime}$ as the same basis. We note, that our results hold also in the general setting where $\mathcal{B}$$^{\prime}$ is any set, not necessarily a basis. In coding theory there are examples where the concept of one-way well-behaving gives better result than the symmetric concepts of well-behaving and weakly well-behaving pairs. We leave it for further research to possibly enhance the theory of the present paper to include structures over skew fields where possibly the one-way feature might be instrumental.
@@
For $\vec{c}=(c_{1},\ldots,c_{n})$ we define ${\mathcal{P}}_{A}(\vec{c})=(d_{1},\ldots,d_{n})$ where $d_{i}=c_{i}$ whenever $i\in A$ and $d_{i}=0$ otherwise from which we define
@@
To ease the exposition we now recall the concept of relative generalized Hamming weights, but here described in the setting of arbitrary field (we already applied such notation in (4)). Given $C_{2}\subsetneq C_{1}\subseteq{\mathbb{F}}^{n}$ write $k_{2}=\dim C_{2}$, $k_{1}=\dim C_{1}$ and $t=k_{1}-k_{2}$. For $1\leq k\leq t$ we define
@@
We are now ready to state our answer to (II) in the setting of linear algebra based on which in the next section we shall prove Theorem 5.
@@
Let $1\leq k_{2}<k_{1}\leq n$ and consider $C_{2}:={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{k_{2}}\}\subseteq C_{1}:={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{k_{1}}\}$. Write $t=k_{1}-k_{2}$ and consider $1\leq k\leq t$. For all $A\subseteq{\mathcal{I}}$ with
$A<\min\{\bar{\mu}(i_{1},\ldots,i_{t-k+1})\mid k_{2}+1\leq i_{1}<\cdots<i_{t-k+1}\leq k_{1}\}$ (16)
@@
We prove that for all $A\subseteq{\mathcal{I}}$ with $A<M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ the result holds and that for some $A$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ it does not. Combining this with (15) (where we substitute $k$ with tk+1t-k+1) finalizes the proof. Given $A$ the number of linearly independent vectors being identically equal to 0 on $A$ is
@@ 4 Proof of Theorem 5
If $\mathbb{F}$} is a finite field, say with $q$ elements, then consider $A_{1}=\cdots=A_{m}={\mathbb{F}}_{q}$. Otherwise, let $q=\mu(M_{1},\ldots,M_{t})$ and choose finite sets $A_{1},\ldots,A_{m}\subseteq{\mathbb{F}}$ all of size at least $q$ in such a way that the $A$ in Theorem 5 satisfies $A\subseteq A_{1}\times\cdots\times A_{m}$. Write $a_{i}=\#A_{i}$, for $i=1,\ldots,m$, $n=a_{1}\cdots a_{m}$ and $\{P_{1},\ldots,P_{n}\}=A_{1}\times\cdots\times A_{m}$. Define $G_{i}:=\prod_{b\in A_{i}}(X_{i}-b)$, for $i=1,\ldots,m$ and consider the ideal $I=\langle G_{1},\ldots,G_{m}\rangle$. The footprint is
@@
and by interpolation the map ${\mbox{ev}}:{\mathbb{F}}[X_{1},\ldots,X_{m}]/I\rightarrow{\mathbb{F}}^{n}$ given by ${\mbox{ev}}(F+I)=(F(P_{1}),\ldots,F(P_{n}))$ is surjective. Preimage and image are of the same size, and therefore $\{{\mbox{ev}}(N)\mid N\in\Delta_{\prec}(I)\}$ is a basis for ${\mathbb{F}}^{n}$ as a vector space over $\mathbb{F}$}. We enumerate the monomials of the footprint according to $\prec$ as $N_{1}\prec\cdots\prec N_{n}$ and write $\vec{b}_{i}={\mbox{ev}}(N_{i})$. From the identity ${\mbox{ev}}(F+I)\ast{\mbox{ev}}(G+I)={\mbox{ev}}(FG+I)$ we immediately see that $\mu(N_{i_{1}},\ldots,N_{i_{k}})\leq\bar{\mu}(i_{1},\ldots,i_{k})$ and therefore the first part of Theorem 5 follows from the first part of Theorem 17. In the remaining part of the proof we assume that $\mathbb{F}$} is the finite field ${\mathbb{F}}_{q}$. Using notation as in [9] consider $W\subseteq\Delta_{\prec}(I)$ and $E(W):={\mbox{Span}}_{\mathbb{F}_{q}}\{{\mbox{ev}}(M)\mid M\in W\}$ and $C(W):=E(W)^{\perp}$. By [9, Pro. 2.4] whenever $W$ is divisor closed it holds that $C(W)=E(W^{\perp})$ where
@@
Using the notation as in Theorem 5 in combination with that of Theorem 17 define
$C_{2}$ $=$ ${\mbox{ev}}\left({\mathcal{L}}_{2}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\prec M_{1}\}\right)$
$C_{1}$ $=$ ${\mbox{ev}}\left({\mathcal{L}}_{1}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\preceq M_{t}\}\right)$
@@
$C_{2}^{\perp}$ $=$ ${\mbox{ev}}\left({\mathcal{L}}_{2}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\preceq\dfrac{N}{M_{1}}\}\right)$
$C_{1}^{\perp}$ $=$ ${\mbox{ev}}\left({\mathcal{L}}_{1}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\prec\dfrac{N}{M_{t}}\}\right)$
where $N=X_{1}^{q-1}\cdots X_{m}^{q-1}$. Clearly ${\mbox{lm}}({\mathcal{L}}^{\perp}_{2}\backslash{\mathcal{L}}_{1}^{\perp})=\left\{\dfrac{N}{M_{t}},\ldots,\dfrac{N}{M_{1}}\right\}$ and the last part of Theorem 5 follows from the last part of Theorem 17 in combination with the last part of Corollary 10.

### v3

@@ 2 Main results on multivariate polynomials
originally introduced for finite fields, but in the present paper applied to the case of any field. Here, ${\mbox{Supp}}(D)$ means the set of entries
$\{i\mid i\in\{1,\ldots,n\},c_{i}\neq 0{\mbox{ for some }}(c_{1},\ldots,c_{n})\in D\}.$
Clearly, for general $F_{1},\ldots,F_{k}$ as in Theorem 5 the smallest attainable number of non-common roots in $A_{1}\times\cdots\times A_{m}$ equals $M_{k}(C_{1},C_{2})$ which as demonstrated in the coming sections equals (5). In the coming sections we shall also demonstrate that if $A\subseteq A_{1}\times\cdots\times A_{m}$ satisfies $A<M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ then there exist $k$ linearly independent polynomials as in Theorem 5 which all evaluates to zero on $A$, but that for some $A\subseteq A_{1}\times\cdots\times A_{m}$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ this does not hold. Here, the dual space (or null-space) is with respect to the usual inner product. Finally, as we shall demonstrate in the next two sections of the paper the right hand side of (1) serves as a lower bound on $M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$. We conclude that problem (I) and problem (II) and our answers to them are basically identical, but applied to dual spaces. All claims of the present remark follows by inspection of the proof of Theorem 5 in Section 4.
We conclude this section by demonstrating the superiority of the footprint bound over other known bounds on roots of general single mulitvariate polynomials. Among these the generalized Alon-Füredi bound [7] is the sharpest as it has as corollary the Alon-Füredi bound [2] which in turn has as corollaries [7] Alon’s combinatorial Nullstellensatz [1] as well as the Schwartz-Zippel bound [47, 41]. For common roots of sets of more than one general multivariate polynomial we are not aware of other efficient bounds than the footprint bound (Theorem 8, Corollary 9 and Corollary 10).
By inspection the generalized Alon-Füredi bound [7, Thm. 1.2] can be reformulated as follows.
@@
In [6] it is already noted that the generalized Alon-Füredi bound is a consequence of the footprint bound. We here give some more details including discussing when the two bounds may or may not coincide. Consider a graded monomial ordering, i.e. a monomial ordering $\prec$ satisfying that if $\deg N_{1}<\deg N_{2}$ then necessarily $N_{1}\prec N_{2}$. There are many such orderings, including $m!$ degree lexicographic ordering where in addition to the above property ties are broken by the rule that if $K_{1}$ and $K_{2}$ are of the same degree then $K_{1}\prec K_{2}$ if $K_{1}\prec_{lex}K_{2}$ for the given chosen lexicographic ordering. For a graded monomial ordering for sure the leading monomial ${\mbox{lm}}(F)$ belongs to the set considered in (7) and Theorem 12 follows imediately from Corollary 10.
We now comment on the possible tightness of the generalized Alon-Füredi bound. If $d_{1}+\cdots+d_{m}=d$ then there is only one monomial in the support ${\mbox{Supp}}(F)$ of highest degree, hence the two bounds coincide. Moreover, $F$ is a monomial ordering invariant polynomial [23, Def. 8], i.e. all monomials in the support of $F$ divides the unique monomial of highest degree. As shown in [23, Cor. 9] the only polynomials that attain the footprint bound in such a case are the products of linear terms as described in the proof of Corollary 10 (see also [7, Sec. 3.3]). Imagine next that $d_{1}+\cdots+d_{m}\neq d$, in which case there need to be more than one monomial in the argument of (7). If for some $i$ no monomial $M$ in the support of $F$ satisfies simultaneously $\deg_{X_{i}}M=d_{i}$ and $\deg M=d$ then for a lexicographic ordering with $X_{j}\prec_{lex}X_{i}$ for all $j\neq i$ the footprint bound will produce a strictly tighter result than Theorem 12. Example 2 below illustrates the situation. Assume on the other hand that for all $i$ such a monomial exists, then a nesecary condition for Theorem 12 to produce as sharp a result as the footprint bound is that for each of the graded monomial orderings the largest monomial in the argument of (7) is of $\sigma$-value equal to (7) which indeed is a very restrictive assumption on the form of $F$.
@@ 3 Results in linear algebra setting
As a preperation for providing in Section 4 a proof of Theorem 5 as well as giving the missing details of Remark 11, in the current section we treat a problem similar to (II), but in the much more general setting of linear codes over arbitrary fields rather than in the context of multivariate polynomials only. To support future applications beyond multivariate polynomials and for the sake of completeness we also include a treatment of a problem similar to (I) in the setting of linear codes. The main result of the section is Theorem 17 which is the linear code version of Theorem 5.
@@
In this section we start by enhancing such results to hold over any field $\mathbb{F}$}, i.e. we demonstrate that the proof of the Feng-Rao bounds holds regardless of underlying field, our exposition closely following that of [3, 28]. With such theory in hand and by applying Forney’s two duality lemmas [21] our results concerning multivariate polynomials will follow as easy corollaries.
@@
${\mbox{Supp}}(D):=\{i\in{\mathcal{I}}\mid c_{i}\neq 0{\mbox{ for some }}(c_{1},\ldots,c_{n})\in D\}$
(we already used this notation in Remark 11). Note, that this is a different support function than the one we have employed for polynomials earlier in the paper. We are interested in establishing information on the size of ${\mbox{Supp}}(D)$ which we call $w_{H}(D):=\#{\mbox{Supp}}(D)$. Here, the subscript $H$ refers to the Hamming weight from coding theory.
Let $B=\{{\vec{b}}_{1},\ldots,{\vec{b}}_{n}\}$ be an ordered basis for ${\mathbb{F}}^{n}$. Define $L_{0}:=\{\vec{0}\}$ and for $i=1,\ldots,n$ $L_{i}:={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{i}\}$. Further define $\bar{\rho}:{\mathbb{F}}^{n}\rightarrow\{0,1,\ldots,n\}$ by $\bar{\rho}(\vec{0}):=0$ and for $\vec{c}\neq\vec{0}$ $\bar{\rho}(\vec{c}):=i$ if $\vec{c}\in L_{i}\backslash L_{i-1}$ and $m(\vec{c}):=\min\{m\mid\vec{c}\cdot\vec{b}_{m}\neq 0\}$. Here, $\cdot$ is the usual inner product.
@@
Consider a subspace $D\subseteq{\mathbb{F}}^{n}$ of dimension $k\geq 1$. There exists a basis $\{\vec{c}_{1},\ldots,\vec{c}_{k}\}$ for $D$ with $\bar{\rho}(\vec{c}_{1}),\ldots,\bar{\rho}(\vec{c}_{k})$ being pairwise different. For such a basis it holds that $m(\vec{c}_{1}),\ldots,m(\vec{c}_{k})$ are also pairwise different. Both the $\bar{\rho}$-values and the $m$-values are invariants of $D$. Hence, if $\vec{c}\in D\backslash\{\vec{0}\}$ then $\bar{\rho}(\vec{c})$ and $m(\vec{c})$, respectively, belongs to the above mentioned sets, respectively.
By Lemma 13 the following definition is clear
$\bar{\rho}(D):=\{e\mid 1\leq e\leq n,\exists\vec{c}\in D\backslash\{\vec{0}\}{\mbox{ with }}\vec{c}\in L_{e}\backslash L_{e-1}\}$
$m(D):=\{m\mid 1\leq m\leq n,\exists\vec{c}\in D\backslash\{\vec{0}\}{\mbox{ with }}\vec{c}\in L^{\perp}_{m-1}\backslash L^{\perp}_{m}\}$
and we have $\bar{\rho}(D)=\#m(D)=k$.
We next introduce the concept of one-way well-behaving pairs which is more general than the well-known concepts of well-behaving pairs and weakly well-behaving pairs [36] allowing for a more powerful description [28]. Consider an additional ordered basis $B^{\prime}=\{\vec{b}^{\prime}_{1},\ldots\vec{b}^{\prime}_{n}\}$ (which may or may not be equal to $B$). Let ${\mathcal{I}}:=\{1,2,\ldots,n\}$. An ordered pair $(i,j)\in{\mathcal{I}}^{2}$ is said to be one-way well-behaving (OWB) if $\bar{\rho}(\vec{b}_{u}\ast\vec{b}^{\prime}_{j})<\bar{\rho}(\vec{b}_{i}\ast\vec{b}^{\prime}_{j})$ for $u<i$. Given bases $B,B^{\prime}$ as above consider for $l=1,2,\ldots,n$ and $i=1,2,\ldots,n$ the following sets
@@
To prove (12) we therefore repeat the proof of [3][Thm. 10] which with our notation goes as follows. Consider a basis $\{\vec{c}_{1},\ldots,\vec{c}_{k}\}$ for $D$ as in Lemma 13. Write $e_{u}:=\bar{\rho}(\vec{c}_{u})$ for $u=1,\ldots k$ and observe that if $(e_{u},j)$ is OWB for some $j\in\{1,\ldots,n\}$ and $\bar{\rho}(\vec{b}_{e_{u}}\ast\vec{b}^{\prime}_{j})=\ell$ then by the very definition of OWB also $\bar{\rho}(\vec{c}_{u}\ast\vec{b}_{j}^{\prime})=\ell$. Hence,
@@
To prove (13) we repeat the proof of the last part of [28][Thm. 1]. Write $\gamma:=\bar{\mu}(m_{1}:=m(\vec{c}_{1}),\ldots,m_{k}:=m(\vec{c}_{t}))$ and
@@
For the purpose of the present paper the reader may think of $\mathcal{B}$} and $\mathcal{B}$$^{\prime}$ as the same basis. We note, that our results hold also in the general setting where $\mathcal{B}$$^{\prime}$ is any set, not necessarily a basis.
@@
For $\vec{c}=(c_{1},\ldots,c_{n})$ we define ${\mathcal{P}}_{A}(\vec{c}):=(d_{1},\ldots,d_{n})$ where $d_{i}:=c_{i}$ whenever $i\in A$ and $d_{i}:=0$ otherwise from which we define
@@
To ease the exposition we now recall the concept of relative generalized Hamming weights, but here described in the setting of arbitrary field (we already applied such notion in (6)). Given $C_{2}\subsetneq C_{1}\subseteq{\mathbb{F}}^{n}$ write $k_{2}=\dim C_{2}$, $k_{1}=\dim C_{1}$ and $t=k_{1}-k_{2}$. For $1\leq k\leq t$ we define
@@
We are now ready to state our answer to (II) in the setting of linear algebra, based on which in the next section we shall deduce Theorem 5.
@@
Let $1\leq k_{2}<k_{1}\leq n$ and consider $C_{2}={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{k_{2}}\}\subseteq C_{1}={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{k_{1}}\}$. Write $t=k_{1}-k_{2}$ and consider $1\leq k\leq t$. For all $A\subseteq{\mathcal{I}}$ with
$A<\min\{\bar{\mu}(i_{1},\ldots,i_{t-k+1})\mid k_{2}<i_{1}<\cdots<i_{t-k+1}\leq k_{1}\}$ (18)
@@
We prove that for all $A\subseteq{\mathcal{I}}$ with $A<M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ the result holds and that for some $A$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ it does not. Combining this with (17) (where we substitute $k$ with tk+1t-k+1) finalizes the proof. Given $A$ one can write the number of linearly independent vectors in $C_{1}\backslash C_{2}$ being identically equal to 0 on $A$ as
@@ 4 Proof of Theorem 5
If $\mathbb{F}$} is a finite field, say with $q$ elements, then consider $A_{1}=\cdots=A_{m}={\mathbb{F}}_{q}$. Otherwise, let $q:=\mu(M_{1},\ldots,M_{t})$ and choose finite sets $A_{1},\ldots,A_{m}\subseteq{\mathbb{F}}$ all of size at least $q$ in such a way that the $A$ in Theorem 5 satisfies $A\subseteq A_{1}\times\cdots\times A_{m}$. Write $a_{i}=\#A_{i}$, for $i=1,\ldots,m$, $n=a_{1}\cdots a_{m}$ and $\{P_{1},\ldots,P_{n}\}=A_{1}\times\cdots\times A_{m}$. Define $G_{i}:=\prod_{b\in A_{i}}(X_{i}-b)$, for $i=1,\ldots,m$ and consider the ideal $I:=\langle G_{1},\ldots,G_{m}\rangle$. The footprint is
@@
and by interpolation the map ${\mbox{ev}}:{\mathbb{F}}[X_{1},\ldots,X_{m}]/I\rightarrow{\mathbb{F}}^{n}$ given by ${\mbox{ev}}(F+I)=(F(P_{1}),\ldots,F(P_{n}))$ is surjective. Preimage and image are of the same size, and therefore $\{{\mbox{ev}}(N+I)\mid N\in\Delta_{\prec}(I)\}$ is a basis for ${\mathbb{F}}^{n}$ as a vector space over $\mathbb{F}$}. We enumerate the monomials of the footprint according to $\prec$ as $N_{1}\prec\cdots\prec N_{n}$ and write $\vec{b}_{i}={\mbox{ev}}(N_{i}+I)$. From the identity ${\mbox{ev}}(F+I)\ast{\mbox{ev}}(G+I)={\mbox{ev}}(FG+I)$ we immediately see that $\mu(N_{i_{1}},\ldots,N_{i_{k}})\leq\bar{\mu}(i_{1},\ldots,i_{k})$ and therefore the first part of Theorem 5 follows from the first part of Theorem 17. In the remaining part of the proof we assume that $\mathbb{F}$} is the finite field ${\mathbb{F}}_{q}$. Using notation as in [9] consider $W\subseteq\Delta_{\prec}(I)$ and $E(W):={\mbox{Span}}_{\mathbb{F}_{q}}\{{\mbox{ev}}(M+I)\mid M\in W\}$ and $C(W):=E(W)^{\perp}$. By [9, Prop. 2.4] whenever $W$ is divisor closed it holds that $C(W)=E(W^{\perp})$ where
@@
Here, by divisor closed we mean that if $M\in W$ and $N\mid M$ then also $N\in W$. Using the notation as in Theorem 5 in combination with that of Theorem 17 define
$C_{2}$ $:=$ ${\mbox{ev}}\left(({\mathcal{L}}_{2}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\prec M_{1}\})+I\right)$
$C_{1}$ $:=$ ${\mbox{ev}}\left(({\mathcal{L}}_{1}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\preceq M_{t}\})+I\right)$
@@
$C_{2}^{\perp}$ $=$ ${\mbox{ev}}\left(({\mathcal{L}}_{2}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\preceq\dfrac{N}{M_{1}}\})+I\right)$
$C_{1}^{\perp}$ $=$ ${\mbox{ev}}\left(({\mathcal{L}}_{1}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\prec\dfrac{N}{M_{t}}\})+I\right)$
where $N:=X_{1}^{q-1}\cdots X_{m}^{q-1}$. Clearly ${\mbox{lm}}({\mathcal{L}}^{\perp}_{2}\backslash{\mathcal{L}}_{1}^{\perp})=\left\{\dfrac{N}{M_{t}},\ldots,\dfrac{N}{M_{1}}\right\}$ and the last part of Theorem 5 follows from the last part of Theorem 17 in combination with the last part of Corollary 10.